# Adversarially-Robust Gossip Algorithms for Approximate Quantile and Mean Computations

### v2

@@ 6       $x_{v}\leftarrow\mathrm{median}(x_{u_{1}(v)},x_{u_{2}(v)},x_{u_{3}(v)})$
Algorithm 1 $3$-TOURNAMENT($v$)
To compute approximate quantiles $\phi$, [HMS18] gives another tournament-style algorithm which reduces the problem to that of finding an approximate median. Our algorithm builds on their ingeniously simple algorithm. Its basic principle can be described in one sentence: In each round, each node pulls two node values uniformly at random and updates its value to be the the minimum (if $\phi<1/2$) or maximum (if $\phi>1/2$) of the two. Our algorithm here requires more refined tracking of its own progress and knowing when to terminate, taking into account the magnitude of the potential corruption, which we measure in terms of $\beta$. This also greatly affects the required analysis. The pseudocode for our quantile shifting algorithm is given in Algorithm 2 (for the case $\phi<1/2$).
@@ 13       $i\leftarrow i+1$
Algorithm 2 $2$-TOURNAMENT($v$)
To compute the approximate mean, [KDG03] gives the elegant push-sum algorithm which roughly works as follows: In each round, every node pushes half of its value to another node chosen uniformly at random. After $O\big{(}\log n+\log\frac{1}{\varepsilon}\big{)}$ rounds, the algorithm converges to an approximate solution. The authors use a crucial property of this push-based protocol – mass preservation – to show that the sample variance of the values decreases exponentially, while the sample mean remains more or less unchanged.
As discussed before, the mass preservation property completely breaks down in the presence of adversarial corruption: all it takes is for a single node to report an abnormally large value once. Moreover, pull-based approaches generally work better than push-based ones in the presence of adversarial faults, as they inherently restrict the influence that adversarial corruptions can have. We adopt a novel approach and design a tournament-style algorithm, but use however a similar proof technique as in [KDG03], despite the absence of mass preservation. The algorithm itself is very simple - each node repeatedly updates its value to the average of two node values chosen uniformly at random. Finally, we also end with a majority vote, as in the approximate median algorithm, to ensure that at most a $\gamma$ fraction of the nodes output a wrong answer. The pseudocode of our adversarially robust mean approximation algorithm can be found in Algorithm 3.
@@ 7       $x_{v}\leftarrow\frac{x_{u_{1}(v)}+x_{u_{2}(v)}}{2}$
We highlight the trade-off between the tournament-style aggregation in the first phase and the sequential collecting of values for the majority vote in the second phase. Though the former is more efficient, it also gives the adversary more (exponentially more in fact) influence over the values.
@@
There are a variety of directions for follow-up work. In terms of other aggregation tasks, beyond computation of median, other quantiles and mean, the main open problem which may robustly be solved in our setting is the computation of the mode (since there is no hope for computing the minimum or maximum value in the presence of adversaries). We also think it would be interesting to investigate the model in the asynchronous setting, or relax the assumption that communications occur on an underlying complete graph.
1.3 Related Work
@@
The Gossip Protocol is also known as the Epidemic Protocol, as these algorithms were first developed to imitate how epidemics spread [DGH+87]. One of the first gossip algorithms to be studied was rumor spreading or randomized broadcast [KSSV00, Pit87], where in [KSSV00] the authors give an algorithm for rumor spreading in $O(\log n)$ rounds and $O(n\log\log n)$ messages.
@@
In [KDG03], the authors study a gossip protocol to compute aggregates such as sums and counts in $O(\log n)$ rounds. Furthermore, they also develop algorithms for random sampling and quantile computation (also known as randomized selection), the latter of which can be computed in $O(\log^{2}n)$ rounds. The work in [HMS18] improves on that and gives an algorithm to compute exact quantiles in $O(\log n)$ rounds and approximate quantiles in $O\big{(}\log\log n+\log\frac{1}{\varepsilon}\big{)}$ rounds. The problem of computing quantiles has also been studied in the centralized setting [BFP+73] and the distributed setting [KLW07]. Other problems such as computing the mode [GP16] (also known as plurality consensus), have also been studied in the gossip setting.
@@
Both in [AGGZ10] and [KSSV00], the authors investigate a gossip-based rumor spreading in the presence of adaptive failures, but the adversary here only has the power to crash certain messages, and cannot send altered messages (which is much more powerful). To the best of our knowledge, this is the first time anyone has investigated a gossip model for general aggregation problems in the presence of adversarial nodes. We note that for the special case of median computation, restricting the adversarial influence to $\beta\leq\frac{1}{n^{1/2}}$, the authors of [DGM+11] derive a runtime bound of $O(\log n\log\log n)$ for the computation of an approximate median from their approach to the stabilizing consensus problem. In this special case, at the cost of an extra factor of $O(\log\log n)$ compared to our runtime, they get an approximation factor of $O\big{(}\frac{(\log n)^{1/2}}{n^{1/2}}\big{)}$, thus yielding a tighter approximation than provided by our Theorem 1, which stipulates that $\varepsilon=\Omega(\frac{1}{n^{0.0019}})$. We note that we did not try to optimize the exponent in this lower bound on $\varepsilon$. In the deterministic setting, approximate median computation in the presence of byzantine nodes has been studied [SW15]. The algorithm presented there was shown to yield an approximation factor of at most $\frac{\beta}{2}$, and runs in time in time $O(\beta n)$, as long as $\beta<\frac{1}{3}$. This latter requirement on the fraction of adversarial nodes was proven to be tight. This was later generalized to approximate $k$-th largest values in the presence of byzantine nodes [MW18].
@@ 2 Approximate Median
The objective of this section is to prove the runtime and correctness guarantees for our adversarially-robust median algorithm, as stated in Theorem 1.
@@
The algorithm, which draws on ideas from [HMS18], is described in Algorithm 1, and proceeds in two phases. In the first phase, every iteration consists of $3$ rounds. More concretely, in every iteration, each node pulls the value of three random nodes, and updates its value to the median of these three values. In the second phase, every node obtains $K$ many independent samples and outputs the median of these $K$ values.
@@
Proof Sketch. We first show that, by the end of the first phase, i.e. after $t$ rounds (line $3$ of the pseudocode of Algorithm 1), $l_{t}$ is small enough (smaller than $\gamma^{\prime}$ as defined on line 2 in Algorithm 1). Next, we show that for all $0\leq i\leq t$, $|L_{i}|/n$ is approximately equal to $l_{i}$ in expectation and with high probability. Thus, after the first phase, at most a $\gamma^{\prime}$ fraction of the nodes have a wrong answer. Finally, we show that after the second phase, this fraction is reduced to $\gamma$.

### v3

@@ 6 $x_{v}\leftarrow\mathrm{median}(x_{u_{1}(v)},x_{u_{2}(v)},x_{u_{3}(v)})$
Algorithm 1 33-TOURNAMENT($v$)
To compute approximate quantiles $\phi$, [HMS18] gives another tournament-style algorithm which reduces the problem to that of finding an approximate median. Our algorithm is based on their ingeniously simple algorithm. Its basic principle can be described in one sentence: In each round, each node pulls two node values uniformly at random and updates its value to be the minimum (if $\phi<1/2$) or maximum (if $\phi>1/2$) of the two. Our algorithm here requires a more refined tracking of its own progress and knowing when to terminate, taking into account the magnitude of the potential corruption, which we measure in terms of $\beta$. This also greatly affects the required analysis. The pseudocode for our quantile shifting algorithm is given in Algorithm 2 (for the case $\phi<1/2$).
@@ 13 $i\leftarrow i+1$
Algorithm 2 22-TOURNAMENT($v$)
To compute the approximate mean, [KDG03] gives the elegant push-sum algorithm, which works roughly as follows: In each round, every node pushes half of its value to another node which is chosen uniformly at random. After $O\big{(}\log n+\log\frac{1}{\varepsilon}\big{)}$ rounds, the algorithm converges to an approximate solution. The authors use a crucial property of this push-based protocol, mass preservation, to show that the sample variance of the values decreases exponentially, while the sample mean remains more or less unchanged.
As discussed before, the mass preservation property breaks down in the presence of adversarial corruption. Additionally, push-based protocols are not ideal as adversarially corrupted nodes gain a significant amount of influence by pushing faulty values to an arbitrary number of nodes (unless we restrict their power). Moreover, pull-based approaches generally work better than push-based ones in the presence of adversarial faults, as they inherently restrict the influence that adversarial corruptions can have. We adopt a novel approach and design a tournament-style algorithm, but use, however, a similar proof technique as in [KDG03], despite the absence of mass preservation. The algorithm itself is very simple - each node repeatedly updates its value to the average of two node values chosen uniformly at random. Finally, we also end with a majority vote, as in the approximate median algorithm, to ensure that at most a $\gamma$ fraction of the nodes output a wrong answer. The pseudocode of our adversarially robust mean approximation algorithm can be found in Algorithm 3.
@@ 7 $x_{v}\leftarrow\frac{x_{u_{1}(v)}+x_{u_{2}(v)}}{2}$
We highlight the trade-off between the tournament-style aggregation in the first phase and the sequential collection of values for the majority vote in the second phase. Though the former is more efficient, it also gives the adversary more (exponentially more in fact) influence over the values.
@@
There are a variety of directions for follow-up work. In terms of other aggregation tasks, beyond computation of median, other quantiles, and mean, the main open problem which may robustly be solved in our setting is the computation of the mode (since there is no hope for computing the minimum or maximum value in the presence of adversaries). We also think it would be interesting to investigate the model in the asynchronous setting or relax the assumption that communications occur on an underlying complete graph.
4 Related Work
@@
The Gossip Protocol is also known as the Epidemic Protocol, as these algorithms were first developed to mimic the spread of epidemics [DGH+87]. One of the first gossip algorithms studied was rumor spreading or randomized broadcast [KSSV00, Pit87], where in [KSSV00] the authors give an algorithm for rumor spreading in $O(\log n)$ rounds and $O(n\log\log n)$ messages.
@@
In [KDG03], the authors study a gossip protocol to compute aggregates such as sums and counts in $O(\log n)$ rounds. Furthermore, they also develop algorithms for random sampling and quantile computation (also known as randomized selection), the latter of which can be computed in $O(\log^{2}n)$ rounds. The work in [HMS18] improves that and gives an algorithm to compute exact quantiles in $O(\log n)$ rounds and approximate quantiles in $O\big{(}\log\log n+\log\frac{1}{\varepsilon}\big{)}$ rounds. The problem of computing quantiles has also been studied in the centralized setting [BFP+73] and the distributed setting [KLW07]. Other problems such as computing the mode [GP16] (also known as plurality consensus), have also been studied in the gossip setting.
@@
In both [AGGZ10] and [KSSV00], the authors investigate a gossip-based rumor spreading in the presence of adaptive failures, but the adversary here only has the power to crash certain messages and cannot send altered messages (which is much more powerful). To the best of our knowledge, this is the first time a gossip model has been investigated for general aggregation problems in the presence of adversarial nodes. We note that for the special case of median computation, restricting the adversarial influence to $\beta\leq\frac{1}{n^{1/2}}$, the authors of [DGM+11] derive a runtime bound of $O(\log n\log\log n)$ for the computation of an approximate median from their approach to the stabilizing consensus problem. In this special case, at the cost of an additional factor of $O(\log\log n)$ compared to our runtime, they get an approximation factor of $O\big{(}\frac{(\log n)^{1/2}}{n^{1/2}}\big{)}$, thus yielding a tighter approximation than provided by our Theorem 1, which stipulates that $\varepsilon=\Omega(\frac{1}{n^{0.0019}})$. We note that we did not try to optimize the exponent in this lower bound on $\varepsilon$. We note that Byzantine-robust gossip algorithms have been widely studied in the context of decentralized machine learning  [HKJ23],  [GDH25a],  [GDH25b].
In the deterministic setting, the approximate median computation in the presence of byzantine nodes has been studied [SW15]. The algorithm presented there was shown to produce an approximation factor of at most $\frac{\beta}{2}$ and runs in time $O(\beta n)$, as long as $\beta<\frac{1}{3}$. The latter requirement on the fraction of adversarial nodes was proven to be tight. This was later generalized to approximate the $k$ -th largest values in the presence of Byzantine nodes [MW18].
@@ 5 Approximate Median
In this section, we sketch the ideas behind our median approximation algorithm and the proof of its correctness and runtime guarantees stated in Theorem 1, whose proof can be found in Appendix B.
@@
The algorithm, which draws on ideas from [HMS18], is described in Algorithm 1, and proceeds in two phases. In the first phase, every iteration consists of $3$ rounds. More concretely, in every iteration, each node pulls the value of three random nodes and updates its value to the median of these three values. In the second phase, every node obtains $K$ many independent samples and outputs the median of these $K$ values.
@@
Proof Sketch. We first show that, at the end of the first phase, that is, after $t$ rounds (line $8$ of the pseudocode of Algorithm 1), $l_{t}$ is small enough (smaller than $\gamma^{\prime}$ as defined in line 2 of Algorithm 1). The proof is a straightforward induction based on the definition of $l_{i}$, but we utilize tailored bounds on the expressions (involving binomial coefficients) that arise from the recursive definition of $l_{i}$.
Next, we show that $\mathbb{E}\left[\frac{|L_{i+1}|}{n}\ \mid\ L_{i}\right]\leq\overline{B}_{3,1}\left(\frac{|L_{i}|}{n}+\beta\right)$. This follows from an analysis of the main loop of the algorithm. Then, using Chernoff’s bounds, we show that $\frac{|L_{i+1}|}{n}$ is concentrated around its expectation.
@@
Notice that $l_{i}$ is defined recursively in such a way as to mirror the evolution of $\frac{|L_{i}|}{n}$ in a single step. In fact, we make this connection more explicit by showing that for all $0\leq i\leq t$, $|L_{i}|/n$ is approximately equal to $l_{i}$ in expectation and with high probability. Thus, we use the bound on $l_{t}$ to conclude that after the first phase, at most a $\gamma^{\prime}$ fraction of the nodes have a wrong answer.
Finally, the second phase amplifies the number of correct nodes in each step and precisely runs the number of steps needed to reduce this fraction of bad nodes from $\gamma^{\prime}$ to $\gamma$.
6 Shifting Quantiles
In this section, we give some intuition underlying our quantile shifting algorithm and the proof of its correctness and runtime guarantees given in Theorem 2, whose proof can be found in Appendix C.
See 2
Note that we can assume $\phi\leq 1/2$ without loss of generality, as the algorithm and analysis are exactly symmetric for the other case.
The pseudocode for our algorithm is described in Algorithm 2. We consider the sets of nodes whose quantiles lie in $\left[0,\phi-\varepsilon\right),\left[\phi-\varepsilon,\phi+\varepsilon\right]$ and $\left(\phi+\varepsilon,1\right]$ at the end of iteration $i$ of the while loop and call these sets $L_{i},M_{i},H_{i}$ respectively. We want to show that $|L_{i}|$ and $|H_{i}|$ decrease rapidly as the algorithm progresses.
$h_{0}\coloneqq h_{0}^{\prime},\qquad h_{i+1}\coloneqq(h_{i}+\beta)^{2},$
Proof Sketch. Let $t$ be the total number of iterations of the while-loop executed by Algorithm 2. We first show that $t$ is bounded by $\mathcal{O}\left(\log\frac{1}{\varepsilon}\right)$ as required. Notice that $h_{i},h_{i}^{\prime}$ are defined so that they start out the same ($h_{0}=h_{0}^{\prime}$) and drift apart very slowly. We make this explicit by showing that after $t$ steps, they are within about a $(1+\varepsilon)$ factor of each other. Next, we show that $\left(\frac{|H_{i}|}{n}-\beta\right)^{2}\leq\mathbb{E}\left[\frac{|H_{i+1}|}{n}\ \mid\ H_{i}\right]\leq\left(\frac{|H_{i}|}{n}+\beta\right)^{2}$. This follows from an analysis of the main loop of the algorithm.
As in the median find algorithm, notice that the recursive definition of $h_{i}^{\prime},h_{i}$ mirrors the evolution of $\frac{|H_{i}|}{n}$. In fact, we make this explicit by giving an upper and lower bound of $\frac{|H_{i}|}{n}$ in terms of $h_{i},h_{i}^{\prime}$, respectively. This is the most technically involved part of the proof, utilizing tedious but straightforward bounds that we tailored for this analysis. Consequently, from the termination condition $h_{t}^{\prime}\leq T\coloneqq\frac{1}{2}-\frac{21\varepsilon}{16}$, we deduce that $\frac{|H_{t}|}{n}$ is around $\frac{1}{2}-\varepsilon$.
The final part of the proof involves showing that $|M_{i}|$ does not change significantly during the entire algorithm and stays around its initial value of $2\varepsilon$. This follows from an analysis of the expected value of $|M_{i+1}|$ given $M_{i}$ and then using Chernoff’s bounds. This would imply that $\frac{|L_{t}|}{n}$ is also around $\frac{1}{2}-\varepsilon$, which concludes our proof.
7 Approximate Mean
In this section, we give a proof sketch for the runtime and correctness guarantees of our adversarially robust approximate mean algorithm stated in Theorem 3. The complete proof can be found in Appendix D.
See 3
The pseudocode is described in Algorithm 3. In the first phase of the algorithm, in each round, every node $u$ samples two nodes and updates its value to their arithmetic mean. Samples of node value that exceed the acceptable value interval $[0,M]$ are cut to the interval limits. In the second phase, the algorithm globally samples a set of nodes uniformly at random and outputs the median value of these nodes. Let $x_{v}(t)$ denote the value $x_{v}$ stored in the node $v$ at the end of iteration $t$. Additionally, let us denote by $V$ the set of all nodes.
To help analyze the progress of the algorithm, we define the potential function.
$\Phi(t)\coloneqq\sum_{u\neq v\in V}(x_{u}(t)-x_{v}(t))^{2},$
where the sum covers all subsets of $V$ of size $2$ (that is, we do not double-count the pairs). Notice that this potential captures the variation in the values stored in the nodes. As the algorithm progresses, we expect the potential to decrease to a small value as the values stored in the nodes converge to the average. We also define the quantity $\psi(t)\coloneqq\sum_{u\in V}x_{u}(t)$ as the sum of the node values. Notice that $\frac{\psi(0)}{n}$ is the true average that we want to approximate.
Proof Sketch. Intuitively, the $2$-tournament algorithm must decrease the variation of the node values in each step as each node updates its value to the average of two other node values. Through a more precise analysis, we show that $\Phi(t)$ approximately halves in expectation in each step. Thus, by using Azuma’s inequality, we can obtain concentration on this expectation and show that at the end of the $T$ steps it has been reduced to less than about a $\beta$ fraction of its original value.
On the other hand, in each step, the influence of the adversary is limited to a $\beta$ fraction of the nodes. This suffices to limit the change in $\psi(t)$ in a single step to about $\beta nM$ in expectation. Using Azuma’s inequality again, we can bound the total change in $\psi(t)$ after $T$ steps by about $\varepsilon nM$.
If the average of the values does not change much but the variance decreases substantially, we can conclude that most of the values must be close to the average. This is exactly what we make next explicit, showing that the number of nodes that can have a wrong answer at the end of the first phase is bounded by a polynomial in $\beta$.
Finally, as in the median find algorithm, the second phase amplifies the number of correct nodes in each step and precisely runs the number of steps needed to reduce this fraction of bad nodes to $\gamma$.
8 Lower Bounds on Round Complexity
In this section, we prove Theorem 4, which yields a general tool for calculating the lower bounds for the round complexity of gossip algorithms in the presence of adversaries. In essence, the theorem quantifies a lower bound on the number of rounds needed for all but a $\gamma$ fraction of the nodes to produce a correct answer, in the presence of $\beta n$ adversaries. Note that a similar lower bound was proven for the non-adversarial setting in [HMS18], but its proof does not extend readily to our case because our parametrized notion of correctness allows a $\gamma$ fraction of nodes to be incorrect in addition to the $\beta n$ adversarial nodes. However, what remains true is that even with the slack provided by $\gamma$, a correct algorithm cannot terminate in fewer than $\Omega\big{(}\frac{\log(1/\gamma)}{\log(1/\beta)}\big{)}$ rounds, since up to this point, more than a $\gamma$ fraction of nodes will have communicated only with bad nodes. Therefore, they will have received no accurate information whatsoever and could not possibly output the correct result. This is formalized in Theorem 4.
See 4
Proof.
First, suppose that $2\gamma>\beta$. Then, notice that $\frac{\log(1/2\gamma)}{\log(1/\beta)}<1$ and hence the claim holds trivially. Thus, we can now assume that $2\gamma\leq\beta$.
Consider a single good node $v$. In one round, the probability of communicating with an adversarially affected node is $\beta$. Thus, after $t<\frac{\log(1/2\gamma)}{\log(1/\beta)}$ rounds, the probability that it has communicated only with adversarial nodes is at least $\left(\beta\right)^{t}>2\gamma$. Thus, in expectation after $t$ rounds, there is at least a $2\gamma$ fraction of such nodes.
Now, notice that the communication partner chosen by an arbitrary node in an arbitrary round is completely independent of what is chosen by any other node in any round. Thus, by the Chernoff bounds, the probability that there is at least a $\gamma$ fraction of the nodes that have only communicated with bad nodes during the first $t<\frac{\log(1/2\gamma)}{\log(1/\beta)}$ rounds is at least
Similarly, one can adapt the proof of [HMS18, Theorem 1.3] to show a $\Omega\left(\log\log\frac{1}{\gamma}+\log\frac{1}{\varepsilon}\right)$ lower bound for these problems. We formalize this statement in Proposition 5 and give its proof in Appendix E.
Proposition 5.
Together, Theorem 4 and Proposition 5 show that our algorithm for the approximate median is optimal. Moreover, our algorithm for the approximate mean is also almost optimal with only an extra $O(\log\frac{1}{\gamma+\beta})$ factor and is optimal whenever $\varepsilon$ is polynomially smaller than $\gamma$ or $\beta$.
Appendix A Tools and Bounds
In this section, we collect useful tools and bounds that we use for our analysis throughout the paper. We begin with a statement of the standard Chernoff bounds.
Lemma 6.
The following simpler statements hold for $0<\delta<1$
Lemma 7.
Suppose that $X_{1},X_{2},\ldots,X_{n}$ are independent indicator random variables, and let $X=\sum_{i=1}^{n}X_{i}$ be their sum. Then, for some $M\geq\mathbb{E}\left[X\right]$ and $0<\delta<1$,
Proof.
Lemma 6 and [HMS18, Lemma A.2] immediately imply the desired claim. ∎
Next, we state Azuma’s inequality.
Lemma 8.
$b_{n,k}(p)=\binom{n}{k}p^{k}(1-p)^{n-k},$
$B_{n,k}(p)=\sum_{0\leq l\leq k}b_{n,l}(p).$
Lemma 9.
Proof.
We have
Thus,
Lemma 10.
Proof.
The claim is clearly true for $\alpha=0$. For $\alpha>0$, note that
$(1-2\alpha)^{n}=1-f(n,\alpha)\cdot\alpha$
where
Next, we prove another upper bound for exponentials.
Lemma 11.
Suppose that $0\leq x\leq M$. Then we have
$e^{\alpha x}\leq 1+\frac{e^{M\alpha}-1}{M}x.$
Proof.
The claim is clearly true for $x=0$. For $x>0$, note that
$e^{\alpha x}=1+f(x)x$
where
$f(x)\coloneqq\frac{e^{\alpha x}-1}{x}$
is monotonically increasing in $x$ in its domain. Thus, we have
$e^{\alpha x}\leq 1+f(M)x\leq 1+\frac{e^{M\alpha}-1}{M}x.$
Lemma 12.
Consider the functions $f(x)\coloneqq\left(\frac{x}{1+\beta}\right)^{2}$ and $g(x)\coloneqq\left(\frac{x+\beta}{1+\beta}\right)^{2}$ for some $0<\beta<1$. Then, for all $\beta^{2}\leq x\leq 1$, we have $f(x)\leq x$ and $g(x)\leq x$.
Proof.
First, consider the function $F(x)\coloneqq f(x)-x$. We have F(0)=f(0)0=0F(0)=f(0)-0=0, and $F((1+\beta)^{2})=f((1+\beta)^{2})-(1+\beta)^{2}=0$. Thus, since $F$ is a quadratic in $x$ with a positive leading coefficient, we must have $F(x)\leq 0$ for all $0\leq x\leq(1+\beta)^{2}$. Consequently, $f(x)\leq x$ for all $\beta^{2}\leq x\leq 1$.
Next, consider the function $G(x)\coloneqq g(x)-x$. We have G(1)=g(1)1=0G(1)=g(1)-1=0, and $G(\beta^{2})=g(\beta^{2})-\beta^{2}=0$. Thus, since $G$ is a quadratic in $x$ with a positive leading coefficient, we must have $G(x)\leq 0$ for all $\beta^{2}\leq x\leq 1$. Consequently, $g(x)\leq x$ for all $\beta^{2}\leq x\leq 1$. ∎
Theorem 13.
$\overline{B}_{(2k+1),k}(p)\leq\frac{1}{2}-\frac{13}{12}\cdot\alpha.$
Proof.
For $l>k+1$, notice that 2l2k1>12l-2k-1>1, and hence we can use Lemma 10 to further bound the expression as
Appendix B Approximate Median
In this section we prove the runtime and correctness guarantees for our adversarially-robust median algorithm, as stated in Theorem 1.
See 1
Remember the following recursive definition:
$l_{0}\coloneqq\frac{1}{2}-\varepsilon,\qquad l_{i+1}\coloneqq\overline{B}_{3,1}(l_{i}+\beta).$